OpenStudy (anonymous):

please help!; how do you find the solutions for x^4-x^2-12=0

OpenStudy (anonymous):

Factor it out!

OpenStudy (anonymous):

(x^2 - 4)(x^2 + 3) = 0

OpenStudy (anonymous):

i did and got 0,3,-4

OpenStudy (anonymous):

i just want to see if im right

OpenStudy (anonymous):

no!

OpenStudy (anonymous):

x^2 - 4 = 0 (x-2)(x+2) = 0 x=2,-2

OpenStudy (cwrw238):

let y = x^2 then y^2 - y -12 = 0 can you solve this?

OpenStudy (anonymous):

wait there are more solutions

OpenStudy (cwrw238):

yup there will be 4 solution because the degree of the equation is 4

OpenStudy (anonymous):

x^2 + 3 = 0 x^2 = -3 \[x = i \sqrt{3}, -i \sqrt{3}\]

OpenStudy (anonymous):

the solutions are: \[x = 2, -2, i \sqrt{3}, -i \sqrt{3}\]

OpenStudy (anonymous):

am I correct @cwrw238

OpenStudy (anonymous):

i understand why there 4 solutions but where did you get them from?

OpenStudy (anonymous):

First we factored them out, didnt we? then we set each factor equal to zero

jimthompson5910 (jim_thompson5910):

It might be easier to let z = x^2 so z^2 = x^4 which allows us to go from x^4-x^2-12=0 to z^2-z-12=0

jimthompson5910 (jim_thompson5910):

solve z^2-z-12=0 and keep in mind that z = x^2 so you can solve for x

OpenStudy (anonymous):

o because it x^2=3 its the square root plus i right

OpenStudy (anonymous):

ok i got it - thank you so much guys

OpenStudy (anonymous):

there you go @jim_thompson5910 :)

OpenStudy (anonymous):

thank you! :)

OpenStudy (anonymous):

this site is really helpful: mathway.com

OpenStudy (cwrw238):

@tegasree - u r correct

OpenStudy (anonymous):

thank you! :)